The line passing through (-2,y) and (1,0) is perpendicular to line joining (3,3) and (-1,1). Find y.
Thank you.

Found 2 solutions by subudear, richwmiller:
Answer by subudear(62)

Let the slope of first line be m1 and slope of second line be m2 then
for both lines to be perpendicular to each other m1*m2 = -1
lets find m1 and m2
m1 = (0 - y) / [1 - (-2)] = -(y)/3
m2 = (1 - 3) / (-1 - 3) = 2/4 = 1/2
m1*m2 = -1
-(y)/3 * 1/2 = -1
-(y)/6 = -1
y = 6

Answer by richwmiller(17219)

find the slope of the line with points (3,3) and (-1,1)
y-y/x-x=m
1-3/-1-3=-2/-4=1/2
the slope of the line perpendicular is -2
y-y/x-x=m
0-y/1-(-2)=-2
-y/3=-2
y=6